Suppose that an airline accepted 12 reservations for a commuter plane with 10 seats.
They know that 7 reservations went to regular commuters who will show up for sure.
The other 5 passengers will show up with a 50% chance, independently of each other.
Find the probability that the flight will be overbooked.
1
Expert's answer
2020-06-02T19:58:14-0400

Find the probability that the flight will be overbooked.

P(X=4)+P(X=5)= C5,4(0.5)4(1−0.5)5−4+C5,5(0.5)5(1−0.5)5−5=(5+1)(0.5)5=0.1875